OpenStudy (anonymous):

Find four consecutive integers whose sums is -38? 1.) -13, -12, -11, -10 2.) -12, -11,-10,-9 3.) -11, -10, -9, -8 4.) -10, -9, -8, -7

OpenStudy (whpalmer4):

We can write 4 consecutive integers as n, n+1, n+2, n+3. We also know that the sum of them = -38. \[n+n+1+n+2+n+3 = -38\] Solve that for n, then compute the next three by n+1, n+2, n+3.

OpenStudy (whpalmer4):

Alternative approach: if you have 4 consecutive integers, you can pair them up as n + n + 3 = 2n + 3 n+1 + n+2 = 2n + 3 Each pair will be half of -38 2n + 3 = -19 I'd do it this way if I didn't have paper and pencil, the other way if I did.

OpenStudy (anonymous):

I reallyyy just need to answer.

OpenStudy (whpalmer4):

You really just need to do the problem. Having the answer without being able to do it is useless.